### v1

# Exploring the ability of the Deep Ritz Method to model strain localization as a sharp discontinuity

Abstract
We present an exploratory study of the possibilities of the Deep Ritz Method (DRM) for the modeling of strain localization in solids as a sharp discontinuity in the displacement field. For this, we use a regularized strong discontinuity kinematics within a variational setting for elastoplastic solids. The corresponding mathematical model is discretized using Artificial Neural Networks (ANNs). The architecture takes care of the kinematics, while the variational statement of the boundary value problem is taken care of by the loss function. The main idea behind this approach is to solve both the equilibrium problem and the location of the localization band by means of trainable parameters in the ANN. As a proof of concept, we show through both 1D and 2D numerical examples that the computational modeling of strain localization for elastoplastic solids within the framework of DRM is feasible.
keywords:
Deep Ritz Method , Physics Informed Neural Networks , Variational methods , Strain localization , Strong discontinuities , Plasticity
[…]
1 Introduction
Strain localization is a material instability that consists in the appearance of zones within a solid in which the strains are highly concentrated. It is a material instability in the sense that it can occur even in regions in which the stress is uniform. Typically, strain localization is related to constitutive models with softening. It is a general phenomenon associated with different types of specific failure mechanisms like fracture in brittle and quasi-brittle materials, or shear bands in geo-materials and metals, which is why its study is of enormous importance from the solid mechanics point of view.
The modeling of strain localization in solids is a very complex task. Standard boundary value problems that model the behavior of solids become ill-posed when strain localization is observed. This is manifested in the tendency of the model to dissipate energy in bands whose width tends to zero. These localized dissipative mechanisms are not easy to capture in standard formulations within the continuum mechanics framework. When the mathematical models are discretized, this ill-posedness is inherited by the numerical model in the form of spurious behavior. Typically, if no regularization is introduced, the dissipation tends to zero as the discretization is refined, which is not what is observed in reality. This lack of objectivity makes the results obtained by the numerical model useless in the post-critical stages.
To solve this lack of objectivity two main strategies have been followed in the literature. On the one hand, for rate-independent models, an intrinsic characteristic length scale is introduced in the constitutive law so that the width of a localization band does not tend to zero. A very common option for this strategy is to use higher-order terms in the governing equations. The intrinsic length scale can also be introduced by adding a scalar field to describe the location of the localization band as a smooth transition between the parts of the material that are localized and those that are not. Then, the band has finite width and the kinematic fields (displacement and strain) do not have discontinuities. In this context, phase fields have been used extensively to simulate fracture with regularized kinematics [1, 2, 3, 4, 5].
On the other hand, there is the strategy of incorporating discontinuities in the kinematics. Strain localization has been approached with methods that use strong discontinuity kinematics in the displacement field (e.g., [6]). Examples of these methods are the strong discontinuity approach (SDA) [7] and the eXtended Finite Element Method (XFEM) [8, 9, 10]. Such approaches have given good results, although they have shown problems when determining both the onset of the localization band and its location.
Regarding the equilibrium of the solid, one particularly powerful option is to use a variational framework. In fact, the most popular phase-field formulations currently used are devised in this framework. Following the work of Francfort and Marigo [11], in which fracture was modeled as a discontinuity in the displacement field, Bourdin et al. [1] proposed a phase-field formulation as an approximation of the problem with a sharp discontinuity. The variational setting has proved to be a rigorous one [12, 1, 13]. In our contribution, we are closer to the original treatment of Francfort and Marigo [11], that is, we consider a variational framework for free discontinuity problems maintaining a kinematics that allows for sharp discontinuities: no localization limiter is introduced in the formulation. In addition, we rely on concepts from the so-called strong discontinuity approach [14] both for the kinematics and to ensure objectivity.
Many options to discretize a boundary value problem (BVP) are available. Perhaps the most popular one in solid mechanics is the Finite Element Method (FEM). Recently, a promising alternative has emerged: the use of artificial neural networks (ANNs) to generate approximation spaces to solve a BVP, the so-called physics-informed neural networks (PINNs) [15]. For instance, it has been recently demonstrated how ReLU-PINNs can emulate the FEM approximation space [16]. He et al. [16] demonstrated how an ANN using Continuous Piece-Wise Linear (CPWL) functions can generate the finite element basis functions. However, they did not present numerical experiments (using Tensorflow or PyTorch, for example).
The variational/energy approach is highly interesting for its implementation using PINNs. A PINN is based on the implementation of a slightly modified ANN. The training mechanics of a PINN is basically the same as that of any ANN: it is based on a network architecture and an optimizer, which uses automatic differentiation to calculate gradients [17]. The fundamental novelty is to consider a loss function containing information of the physical phenomena [18, 19]. In general, the loss function is formulated based on the residuals of the governing equations. However, for some mechanical problems, it is more natural to use the variational approach [12] involving energy minimization. This approach was presented by Weinan and Yu [20] as the Deep Ritz Method (DRM). Within this framework, Samaniego et al. [21] presented examples for many problems in solid mechanics including the phase-field modeling of fracture. Further contributions to the phase-field modeling of fracture can be found, for instance, in [22] and, more recently, [23].
Attempts to model strain localization as as a sharp discontinuity using the variational statement of the BVP and discretized by means of Neural Networks (NNs) are not found in the literature, perhaps with the only exception of [24]. However, sensu stricto, this very interesting proposal uses a Reproducing Kernel Particle Method enhanced with an NN. Here, we explore the possibility of modeling strain localization in elastoplastic solids within the DRM framework, using NNs as the only approximation space for the displacement field, which includes (regularized) discontinuities. On the one hand, the strong discontinuity kinematics is taken care of by the architecture. On the other hand, the variational statement of the boundary value problem is taken care of by the loss function. The main idea behind this approach is to solve both the equilibrium problem and the location of the localization band by means of trainable parameters in the ANN. It might be said that, in a sense, we are solving numerically a free discontinuity problem without resorting to the introduction of regularizing gradients or the addition of a new field (e.g., a phase field). We show, as a proof of concept, that the computational modeling of strain localization as a sharp discontinuity for elastoplastic solids within the framework of DRM is feasible for both 1D and 2D numerical examples.
2 A free discontinuity problem
From a macroscopic point of view, the localization band can be described as a jump discontinuity surface of the displacement field, which is called in the computational mechanics community a strong discontinuity [25].
Within the framework of the so-called free discontinuity problems [26], this kinematics is related to an energy functional $\mathcal{W}$ that can be additively decomposed into a bulk energy $\mathcal{E}$ and a surface energy $\mathcal{S}$ [12]:
$\mathcal{W}[\boldsymbol{u},\Gamma]=\mathcal{E}[\boldsymbol{u},\Gamma]+\mathcal{S}[\llbracket\boldsymbol{u}\rrbracket,\Gamma],$ (1)
where $\llbracket\boldsymbol{u}\rrbracket$ is the jump in displacement field $\boldsymbol{u}$ at the discontinuity surface $\Gamma$. In free discontinuity problems, both $\boldsymbol{u}$ and $\Gamma$ are unknowns.
Here we consider an approximation of this surface by a band of constant width $h$ centered at $\Gamma$:
$\Gamma_{h}\approx\Gamma.$ (2)
This implies considering a regularization of the strong discontinuity kinematics. It is important to mention that $h$ is, in principle, arbitrarily small and that, when $h$ goes to zero, the jump discontinuity at $\Gamma$ is recovered.
3 Kinematics of regularized strong discontinuities
A strong discontinuity can be regularized in the form of a so-called weak discontinuity, where the width of the localization band is no longer equal to zero, but a small positive parameter $h$. Consider the following additive decomposition of the displacement field, $\boldsymbol{u}$ (Figure 1):
$\boldsymbol{u}=\bar{\boldsymbol{u}}+\mathcal{H}_{\Gamma_{h}}\llbracket\boldsymbol{u}\rrbracket,$ (3)
where $\bar{\boldsymbol{u}}$ is the continuous part of the displacement, $\mathcal{H}_{\Gamma_{h}}$ is a regularized Heaviside that varies linearly across $\Gamma_{h}$, and $\llbracket\boldsymbol{u}\rrbracket$ is the jump. From this decomposition and defining the strain as the symmetric gradient of the displacement, one obtains the additive decomposition of the strain tensor
$\boldsymbol{\epsilon}=\nabla^{s}\boldsymbol{u}=\hat{\boldsymbol{\epsilon}}+\frac{1}{h}\mu_{\Gamma_{h}}\llbracket\boldsymbol{u}\rrbracket\odot\boldsymbol{n},$ (4)
where $\hat{\boldsymbol{\epsilon}}$ is the regular part of the strain, $\mu_{\Gamma_{h}}$ is a collocation function acting in $\Gamma_{h}$, and $\boldsymbol{n}$ is the unit normal to the discontinuity surface $\Gamma$.
Figure 1: Kinematic decomposition for a one-dimensional problem. Strong discontinuity is recovered as $h$ tends to zero.
To model plasticity, we consider the additive decomposition of the strain:
$\boldsymbol{\epsilon}=\boldsymbol{\epsilon}^{e}+\boldsymbol{\epsilon}^{p}.$ (5)
We also define the equivalent plastic strain $p$ as usual:
[…]
Notice that
$\dot{\boldsymbol{\epsilon}}^{p}=\dot{p}{\boldsymbol{m}},$ (7)
where $\boldsymbol{m}$ is the so-called plastic flow direction.
For simplicity, we will assume here that
$\dot{\boldsymbol{\epsilon}}^{p}=\frac{1}{h}\mu_{\Gamma_{h}}\llbracket\dot{\boldsymbol{u}}\rrbracket\odot\boldsymbol{n}.$ (8)
This implies that plastic flow and localized strains co-evolve within $\Gamma_{h}$, while the regular strains in the bulk evolve elastically. In addition, if one assumes that $p=0$ at $t=0$ and that plasticity, once initiated, evolves monotonically without elastic unloading, then
$p=\frac{1}{h}\mu_{\Gamma_{h}}\|{\llbracket\boldsymbol{u}\rrbracket}\|.$ (9)
As a consequence,
$\boldsymbol{\epsilon}^{p}=p{\boldsymbol{m}}=\frac{1}{h}\mu_{\Gamma_{h}}\llbracket\boldsymbol{u}\rrbracket\odot\boldsymbol{n}.$ (10)
We will also consider that plastic flow is isochoric:
$\operatorname{Tr}(\boldsymbol{\epsilon}^{p})=\frac{1}{h}\mu_{\Gamma_{h}}\llbracket\boldsymbol{u}\rrbracket\cdot\boldsymbol{n}=0.$ (11)
Notice that this implies that the jump is orthogonal to the unit normal to $\Gamma$, as in mode II fracture.
4 Variational modeling of elasto-plastic solids
We consider the following functional for an elastoplastic solid occupying a volume $\Omega$:
[…]
where $\Psi^{e}$ is the elastic energy density, $H$ is the softening parameter, and $\sigma_{p}$ is the yield stress.
From Equation (9), the energy functional can be written as
$\mathcal{W}_{h}[\boldsymbol{u},p]=$ $\int_{\Omega}\Psi^{e}(\boldsymbol{\epsilon}^{e}(\boldsymbol{u},\boldsymbol{\epsilon}^{p}))dV+\int_{\Gamma_{h}}\bigg{(}\frac{1}{2h}\bar{H}\|\llbracket\boldsymbol{u}\rrbracket\|^{2}+\frac{1}{h}\sigma_{p}\|{\llbracket\boldsymbol{u}\rrbracket}\|\bigg{)}\,dV,$ (13)
where $\bar{H}$ is given by
$H=h\bar{H}.$ (14)
The parameter $\bar{H}$ is termed intrinsic softening parameter and is considered a material property. Notice that, when the bandwidth $h$ tends to zero, we obtain
$\mathcal{W}[\boldsymbol{u},p]=$ $\int_{\Omega}\Psi^{e}(\boldsymbol{\epsilon}^{e}(\boldsymbol{u},\boldsymbol{\epsilon}^{p}))dV+\int_{\Gamma}\bigg{(}\frac{1}{2}\bar{H}\|\llbracket\boldsymbol{u}\rrbracket\|^{2}+\sigma_{p}\|{\llbracket\boldsymbol{u}\rrbracket}\|\bigg{)}\,dS.$ (15)
As a consequence, the last two terms of Equation (15) tend to surface energies on $\Gamma$. From this, we can see that a cohesive energy density is induced:
$\psi(\llbracket\boldsymbol{u}\rrbracket)=\frac{1}{2}\bar{H}\|\llbracket\boldsymbol{u}\rrbracket\|^{2}+\sigma_{p}\|{\llbracket\boldsymbol{u}\rrbracket}\|.$ (16)
The corresponding cohesive force (per unit surface) reads
$\boldsymbol{t}_{c}(\llbracket\boldsymbol{u}\rrbracket)=\sigma_{p}\,\boldsymbol{e}_{T}+\bar{H}\llbracket\boldsymbol{u}\rrbracket,$ (17)
where $\boldsymbol{e}_{T}$ is a unit vector in the direction of $\llbracket\boldsymbol{u}\rrbracket$.
5 Implementation: architectures and training
We propose two architectures that are devised in order to approximate the regularized strong discontinuity kinematics presented above. The following criteria are used:
  1. 1.
    An ANN that captures the regular behavior of the displacement field. This NN must be expressive enough to capture the behaviour of the solid in the elastic domain. For the 1D case, we use a shallow ReLU-PINN to emulate a FEM approximation space. Meanwhile, a multilayer perceptron architecture with 4 hidden layers, each containing 10 neurons, is used for the 2D example, with the ReLU$(\mathbf{x})$ as the activation function.
  2. 2.
    An NN devised to capture the discontinuous part of the displacement field through a customized activation function $\Phi$.
Thus, the NN approximation of the displacement field is constructed as
$\begin{split}\boldsymbol{u}\approx\boldsymbol{u}^{*}&=\boldsymbol{u}_{R}+\boldsymbol{u}_{J},\\ \bar{\boldsymbol{u}}\approx\boldsymbol{u}_{R}&=NN(\mathbf{x}),\\ \mathcal{H}_{\Gamma_{h}}\llbracket\boldsymbol{u}\rrbracket\approx\boldsymbol{u}_{J}&=\mathbf{w}\cdot\Phi(\mathbf{x}),\\ \end{split}$ (18)
where $\boldsymbol{u}_{R}$ is the approximation of the continuous part of the displacement field and $\boldsymbol{u}_{J}$ is the approximation of the jump part. The field $\boldsymbol{u}_{J}$ includes the customized activation function $\Phi$. The vector $\mathbf{w}$ captures the jump on $\Gamma_{h}$ which, due to isochoricity of the plastic strain, develops in the direction tangential to $\Gamma$. The values of $\Phi$ are constrained between 0 and 1.
The activation function $\Phi$ (Figure 2) is defined by
$\begin{split}y&=\boldsymbol{x}\cdot\boldsymbol{n},\\ z&=\frac{y-y_{p}}{c},\\ \Phi&=S(z+0.5;\beta)-S(z-0.5;\beta),\end{split}$ (19)
where $\boldsymbol{x}$ is the position vector, $\boldsymbol{n}$ is the unit vector normal to $\Gamma$, a trainable parameter, $y_{p}$ is also a trainable parameter that controls the position of the midpoint of the localization band, $c$ is a parameter that controls the width of the band, $\beta$ is a parameter that controls the sharpness of the band, and $S$ is the softplus function defined as $S(z,\beta)=\frac{1}{\beta}\ln{(1+e^{\beta z})}$.
Figure 2: Activation function. Left: effect of varying $\beta$ on $\Phi$. Right: effect of varying $h$ on $\Phi$, with $\beta$ fixed at 100.
Algorithm 1 Physics-Informed Neural Network (PINN) Training
1:Input: Physics model (Variational formulation), boundary conditions
2:Output: Displacement field $\boldsymbol{u^{*}}$
3:Initialize neural network parameters $\boldsymbol{\theta}$ with the elastic state
[…]
5:Define the loss function $\mathcal{L}(\boldsymbol{\theta})$:
  • 1.
    $\mathcal{L}_{\text{VF}}(\boldsymbol{\theta})=\mathcal{W}[\boldsymbol{u^{*}}(\boldsymbol{\theta}),p(\boldsymbol{\theta})]$triangleright Variational Formulation loss
  • 2.
    $\mathcal{L}_{\text{BC}}(\boldsymbol{\boldsymbol{\theta})}=\frac{1}{N_{bc}}\sum_{j=1}^{N_{bc}}\left\|\boldsymbol{u}^{*}(\boldsymbol{x}_{j};\boldsymbol{\theta})-\boldsymbol{u}_{bc}(\boldsymbol{x}_{j})\right\|^{2}$triangleright Boundary condition loss
  • 3.
    if $\mathcal{L}_{\text{BC}}(\boldsymbol{\theta})>\lambda$ then
    • $\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}_{\text{BC}}(\boldsymbol{\theta})$triangleright Loss considering only the boundary conditions
  • 4.
    else
    • $\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}_{\text{VF}}(\boldsymbol{\theta})+\mathcal{L}_{\text{BC}}(\boldsymbol{\theta})$triangleright Total loss including energy of the solid
6:while not converged do
  • 1.
    Compute the gradients of $\mathcal{L}(\boldsymbol{\theta})$
  • 2.
    Update $\boldsymbol{\theta}$ using AdamW optimizer
7:end while
8:Redefine the loss function $\mathcal{L}(\boldsymbol{\theta})$: $\triangleright$ Done for 1D only 1. $\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}_{\text{VF}}(\boldsymbol{\theta})+\mathcal{L}_{\text{BC}}(\boldsymbol{\theta})$triangleright Loss considering both the energy and BC loss
9:while not converged do $\triangleright$ Done for 1D only 1. Compute the gradients of $\mathcal{L}(\boldsymbol{\theta})$ 2. Update $\boldsymbol{\theta}$ using L-BFGS optimizer
10:end while
11:Return the trained neural network model
Regarding the training process, the following steps are followed (algorithm 1):
  1. 1.
    The parameters of the neural network are initialized starting from the last elastic state, i.e., the last possible displacement without localization.
  2. 2.
    The energy of the solid is computed from Equation (12) and also, for the loss function, the error of the boundary conditions is considered.
  3. 3.
    If the error of the boundary conditions is greater than a certain threshold $\lambda$, the loss function is calculated only considering this error; otherwise, the loss function incorporates the energy of the solid.
  4. 4.
    The loss function for the 1D example is minimized using the AdamW optimizer for 3000 epochs, followed by a refinement with the L-BFGS optimizer. For the 2D example, the loss function is minimized using the AdamW optimizer for 5000 epochs.
In order to optimally explore the solution space of the NN, a learning rate schedule is used to adapt the way the AdamW optimizer varies at each step: at the beginning of the process, there are larger “steps” to avoid possible local minima or saddle points, where the loss function might get stuck. In this paper we use the CosineDecay() learning schedule, implemented in TensorFlow.
6 Numerical experiments
In this section, we present results both for 1D and 2D cases. The fundamental idea is to show that the proposed methodology is able to accurately reproduce the solution of the variational problem. Given that the examples provided can be solved analytically, we are able to show that the NN solution is, modulo some numerical errors, the solution of the original mathematical problem. This shows that the approximation space given by the proposed architectures contains the exact solution for the cases considered here, which suggests that appropriate approximation spaces can be devised within the framework of PINNs for the solution of strain localization problems.
The section is organized as follows. The results for the displacement field are shown for several loading stages, both for 1D and 2D examples. Then, for each case, the details of how the solution was obtained are analyzed and discussed.
6.1 1D Example
The first example consists of a one-dimensional bar whose domain is the open interval $\Omega=(0,L)$, where $L=10$ is the length of the bar. The bar is fixed at one end ($x=0$), while, at the other end ($x=L$), the displacement is equal to $\delta$ (i.e., $u(L)=\delta$). The bar has a variable cross-sectional area in order for the problem of determining the location of the localization band to be uniquely defined. Specifically, the cross-sectional area $A(x)$ has a parabolic variation, with a minimum value of 1 at the midpoint and maximum values of 2 at the ends, as shown in Figure 3. Then, considering homogeneous material properties, the band is expected to localize at the middle of the bar.
Figure 3: Geometry configuration of the 1D problem.
Figure 4: Evolution of the displacement field for increasing prescribed displacement, showing the NN-predicted solution (solid green lines) and the expected analytical solution (dotted blue lines).
Figure 5: Induced cohesive energy and cohesive force, showing the NN-predicted solution (solid green lines) and the expected analytical solution (dotted blue lines).
Figure 6: Stress and force along the bar as the loading process evolves.
The material of the bar is characterized by the following parameters: $\sigma_{p}=1$, $E=2$ and $\bar{H}=-2/11$. Notice that, in the regularized strong discontinuity approach used here, the bandwidth $h$ is not a material parameter given that, in principle, $h$ can take any sufficiently small value, and the plastic evolution is uniquely determined by the intrinsic softening modulus $\bar{H}$.
In this example, we use 1001 collocation points uniformly distributed along the bar domain. For the regular part of the displacement field, the NN used emulates an FEM approximation space with 11 equally spaced nodes. The parameters that describe the position of the band and the magnitude of the jump were initialized at 0.4 and 0.0, respectively.
Figure 4 shows that the predicted displacement field for various prescribed displacements $\delta$ closely matches the analytical solution. The regular part of the displacement has been approximated using 11 ReLU-like nodes, which amount to a piecewise linear approximation. The NN also demonstrated a strong capability in approximating the jump in the displacement field. For all imposed displacements $\delta$, the training process led to accurate estimation of the jump magnitude within the localization band. In addition, the parameter describing the location of the band was effectively learned, with a noticeable difference with respect to the analytical solution occurring only for $\delta=5.5$. This shows the presence of the solution within the loss function landscape.
Figure 7: Displacement, strain, stress and force along the bar as the loading process evolves, showing the NN-predicted solution (solid green lines) and the expected analytical solution (dotted blue lines).
Figure 8: Stress and force along the bar as the training process evolves.
Figure 5 shows that the induced cohesive energy and force computed by the NN closely matches the analytical solution corresponding to the strong discontinuity approach. This shows that the regularized strong discontinuity kinematics and the variational statement of the BVP are able to capture the appearance of a cohesive band once softening triggers localization.
Figure 6 illustrates the evolution of stress and force for increasing displacements. It is important to note that the stress field remains continuous even in the presence of strain localization within the band. Also, the force remains constant over the entire domain, as expected. For the case of $\delta=5.5$, where the loading process ends, both the stress and force approximate their actual value of 0, despite the minor error in the training of the band’s position.
Here henceforth, we present an analysis using the specific case of $\delta=4.5$. Figure 7 shows the displacement, strain, stress and force fields for this stage. This figure shows that the parameters describing the localization band were correctly learned. In Figure 7 (b), both the location of the band and the magnitude of the jump are closely aligned with the analytical solution for the regularized strong discontinuity problem.
The training process of the displacement field and the jump part of the displacement field are shown in Figure 8 (a-c) and Figure 8 (d-f), respectively. The latter is particularly important as it reveals how the localization band is trained using the AdamW and L-BFGS optimizers. The first 3000 epochs of training of the jump using the AdamW optimizer are plotted in Figure 8 (d,e). There, we observe that the parameter describing the band’s position is the first to be learned. In contrast, the magnitude of the jump continues to be trained. Then, during the training of the band using the L-BFGS optimizer, illustrated in Figure 8 (f), the magnitude of the jump is finally learned. In fact, for all other cases of $\delta$ in the one-dimensional example, a similar pattern was observed: the position of the localization band was primarily learned using the AdamW optimizer, while the magnitude of the jump continued to be trained using AdamW even when the band’s position has almost reached its actual value, reaching its correct value with L-BFGS optimization.
6.2 2D Example
The 2D case consists of a problem involving a square specimen of unit length subjected to simple shear. The specimen is subjected to Dirichlet boundary conditions, both at the bottom side, where it is fixed, and at the top side, where a tangential displacement of magnitude $\delta$ is applied.
The material parameters are $E=5.6$, $\nu=0.4$, $\bar{H}=-1$, and $\sigma_{p}$, with the latter varying across the domain. Specifically, the yield stress $\sigma_{p}$ varies parabolically, with a minimum value of 0.75 along the horizontal line at the vertical center of the domain, and with a maximum value of 1 along both the top and bottom surfaces. This ensures that the localization band forms with a uniquely defined location and orientation.
We use three different bandwidth values, $h\in\{0.05,0.1,0.2\}$, to show the objectivity of the solution with respect to the regularization parameter and, hence, with respect to the width of the localization band. In all cases, the domain contains a grid of 101 $\times$ 101 uniformly distributed collocation points along the $x$ and $y$ axis.
The parameters describing the localization band are initialized as follows: the location of the band $y_{p}$ is set to 0.25; the orientation of the band is set to an angle of $\pi/2$ rads clockwise with respect to the horizontal axis; and the jump magnitude is set to zero for both the horizontal and vertical components of the displacement field.
Figure 9: Evolution of the displacement field with $h=0.2$ for increasing prescribed displacements.
Figure 10: Evolution of the displacement field with $h=0.1$ for increasing prescribed displacements.
Figure 11: Evolution of the displacement field with $h=0.05$ for increasing prescribed displacements.
Figure 12: Induced cohesive energy and cohesive force for different bandwidth values $h$ (solid lines), highlighting the objectivity of the solution and comparing the results with the expected analytical solution (dotted blue lines).
Figure 13: Localized shear strain field and the corresponding uniform shear stress field.
Figure 14: Shear strain field as the training process evolves.
Figures 9, 10 and 11 show the deformed specimen for different prescribed displacements $\delta$, considering different bandwidths $h$ for each figure. We observe in all cases homogeneous strains at $\delta=0.375$ and localization triggered at $\delta=0.4$. As expected, the shear strains grow as the localization band shrinks with $h$, while the displacement magnitude remains invariant. Furthermore, notably, Figure 12 illustrates that the present framework yields objective solutions, insensitive to the choice of bandwidth. In particular, we show that the cohesive energy and cohesive force are uniquely determined by the intrinsic softening parameter, as expected from Equations (16) and (17). Moreover, the NN-predicted values closely match the expected analytical solution.
Additionally, we show that the method is able to capture the uniformity of the stress, although the strain is concentrated in the band. This result can be observed in Figure 13. Figure 14 further illustrates the training process. It is especially remarkable that both the direction and the position of the band are captured by means of trainable parameters.
7 Conclusions
We have used 1D and 2D numerical examples as proof of concept that plastic solids undergoing strain localization can be, in principle, modeled using the DRM combined with regularized strong discontinuity kinematics to approximate the displacement field. Furthermore, this result suggests that variational versions of PINNs hold promise for solving complex computational solid mechanics problems, such as materials with strain localization.
The present study sets the stage for subsequent developments on several fronts. The idea of describing strong discontinuities via trainable ANNs can be applied to many cases beyond softening plasticity, e.g., frictional-dilatant plasticity or damage, and with more general loading conditions. Moreover, while we have shown feasibility for the onset and evolution of fixed localization bands, applying the proposed scheme to general boundary value problems requires handling crack/band propagation. A possible way forward is to develop multiscale techniques with ANN-based discontinuities propagating through a macroscopic mesh.
Acknowledgements
A. V.-P. acknowledges to the VIUC for supporting him through “Conjunto de horas 1”.